A steam turbine receives a steam flow of 1.35 kg/s and delivers 500 kW. The heat loss fromthe casing is negligible. a) Find the change in specific enthalpy across the turbine when the velocities at entrance andexit and the differences in elevation at entrance and exit are negligible.b) Find the change in specific enthalpy across the turbine when the velocity at entrance is 60m/s, the velocity at exit is 360 m/s and the inlet pipe is 3 m above the exhaust pipe
1
Expert's answer
2019-03-15T12:31:56-0400

a) v1 = v2, Z1 = Z2, Q = 0

Find the work done

W = P/m

W = 500 kW / 1.35 kg/s = 370 kJ/kg

Since ∆KE = 0 and ∆PE = 0

We get

∆h = W

∆h = 370 kJ/kg


b) v1 = 60 m/s, v2 = 360 m/s, Z1 = 0, Z2 = 3 m, Q = 0

W= ∆h + ∆KE + ∆PE

∆KE = (v22 - v12) / 2 = ((360 m/s)2 – (60 m/s)2) / 2 = 63 kJ/kg

∆PE = g × (Z2 - Z1) = 9.81 × 3 = 29.43 J/kg = 29.43 ×10-3 kJ/kg

We find ∆h

∆h = W – (∆KE + ∆PE) 

Finally

∆h = 370 kJ/kg – (63 kJ/kg + 29.43 ×10-3 kJ/kg) = 307 kJ/kg